# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ 3.2. Almost para-CR structure on the correspondence space
Given a projective structure []delimited-[][$\nabla$][ ∇ ] on a 3-manifold $M,$ the projectivized cotangent bundle, $\mathbb{P}T^{*}M,$ is equipped with an almost para-CR structure $\mathscr{F}=\mathscr{V}\oplus\mathscr{E}\subset T\mathbb{P}T^{*}M$ where $\mathscr{F}$ is the canonical contact distribution on $\mathbb{P}T^{*}M,$ $\mathscr{V}$ is the vertical distribution for the projection $\mathbb{P}T^{*}M\to M$ and $\mathscr{E}$ is obtained from the horizontal distributions defined by connections in [].delimited-[][$\nabla$].[ ∇ ] .
@@
Recall that $(2n-1)$-dimensional para-CR geometries are modeled by the flag variety of pairs of incident lines $\ell$ and hyperplanes $H$, i.e.
(3.5) $G/P_{1,n}\cong\mathbb{F}_{1,n}(\mathbb{R}^{n+1}):=\{(\ell,H)\ \vline\ \ell\subset H\subset\mathbb{R}^{n+1}\}\subset\mathbb{P}^{n}\times(\mathbb{P}^{n})^{*},$
where $G=\mathrm{SL}(n+1,\mathbb{R})$ and $P_{1,n}$ is the parabolic subgroup that stabilizes a chosen origin. The homogeneous space $\mathbb{F}_{1,n}(\mathbb{R}^{n+1})$ can be identified with $\mathbb{P}T^{*}\mathbb{P}^{n},$ i.e. the projectivized cotangent bundle of the projective space $\mathbb{P}^{n},$ equipped with its canonical contact structure and Legendrian foliations given by the fibers of the natural fibrations $\mathbb{P}T^{*}\mathbb{P}^{n}\to\mathbb{P}^{n}$ and $\mathbb{P}T^{*}\mathbb{P}^{n}\to(\mathbb{P}^{n})^{*}$.
Thus, almost para-CR structures in dimension 5 are Cartan geometries of type $(\mathrm{SL}(4,\mathbb{R}),P_{1,3})$ and are among the so-called regular and normal parabolic geometries, where $P_{1,3}=P_{0}\ltimes P_{+}$ in which $P_{+}$ is the 5-dimensional nilradical and $P_{0}=\mathbb{R}^{\times}\times\mathrm{GL}(2,\mathbb{R})\subset\mathrm{GL}(3,\mathbb{R}).$ In the case of oriented almost para-CR structure on an oriented projective structure induced on the oriented projectivized cotangent bundle, one has $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R}).$ The harmonic curvature of 5-dimensional almost para-CR structures has three components, also known as the fundamental invariants, and can be written as
@@
The vanishing of $K_{H}$ implies local flatness, i.e. local equivalence to the flat model described previously. The vanishing of torsion components $T^{1}:\bigwedge\nolimits^{2}\mathscr{E}\to\mathscr{V}$ and $T^{2}:\bigwedge\nolimits^{2}\mathscr{V}\to\mathscr{E}$ is equivalent to the integrability of the rank 2 distributions $\mathscr{E}$ and $\mathscr{V},$ respectively. Moreover, the vanishing of $W$ is equivalent to the existence of a 4-parameter family of Legendrian surfaces that are null with respect to the indefinite subconformal structure on the contact distribution, e.g. see [KM24] for a review. An almost para-CR structure of dimension $12n-$ 1 arises from a projective structure on an $n$-dimensional manifold in the manner described above, if and only if $W=0$ and $T^{2}=0,$ e.g. see [Čap05],
@@
(3.6) $\varphi=\begin{pmatrix}-\eta^{i}_{i}&\mu_{2}&\mu_{1}&\mu_{0}\\ \tau^{2}&\eta^{2}_{2}&\eta^{2}_{1}&\mu_{4}\\ \tau^{1}&\eta^{1}_{2}&\eta^{1}_{1}&\mu_{3}\\ \tau^{0}&\tau^{4}&\tau^{3}&\eta^{0}_{0}\end{pmatrix},\quad\boldsymbol{\varphi}=\mathrm{d}\varphi+\varphi{\,{\wedge}\;}\varphi=\begin{pmatrix}0&\boldsymbol{\mu}_{2}&\boldsymbol{\mu}_{1}&\boldsymbol{\mu}_{0}\\ 0&\boldsymbol{\eta}^{2}_{2}&\boldsymbol{\eta}^{2}_{1}&\boldsymbol{\mu}_{4}\\ 0&\boldsymbol{\eta}^{1}_{2}&\boldsymbol{\eta}^{1}_{1}&\boldsymbol{\mu}_{3}\\ 0&\boldsymbol{\tau}^{4}&\boldsymbol{\tau}^{3}&\boldsymbol{\eta}^{0}_{0}\end{pmatrix},$
@@
where, using (3.3) and working modulo $\{\tau^{0}\}$, the 2-forms entries $\boldsymbol{\tau}^{3}$ and $\boldsymbol{\tau}^{4}$ in $\boldsymbol{\varphi}$ satisfy
(3.8) $\boldsymbol{\tau}^{3}\equiv W^{0}_{112}\tau^{1}{\,{\wedge}\;}\tau^{2},\quad\boldsymbol{\tau}^{4}\equiv W^{0}_{212}\tau^{1}{\,{\wedge}\;}\tau^{2}.$
@@
(3.9) $\mathscr{E}=\langle\partial_{p^{1}},\partial_{p^{2}}\rangle,\qquad\mathscr{V}=\langle v_{1},v_{2}\rangle,$
@@ 3.3. The twistor bundle and its 2-nondegenerate CR structure
Given an oriented projective structure []delimited-[][$\nabla$][ ∇ ] on $M,$ let $(\mathbb{P}_{+}T^{*}M;\mathscr{E},\mathscr{V})$ be its corresponding almost para-CR structure on the oriented projectivized cotangent bundle, $\mathbb{P}_{+}T^{*}M,$ with co-oriented contact distribution $\mathscr{F}=\mathscr{E}\oplus\mathscr{V}$. We define the twistor bundle of []delimited-[][$\nabla$][ ∇ ], denoted by $\nu\colon T^{[\nabla]}\to\mathbb{P}_{+}T^{*}M,$ as the bundle of almost complex structures on $\mathscr{F}$ that are compatible with its conformal symplectic 2-form and indefinite bilinear form. As a result, at every $p\in\mathbb{P}_{+}T^{*}M,$ any such compatible almost complex structure acts on $\mathscr{F}_{p}$ via the action of the structure group. Recall from § 3.2 that almost para-CR structures on $\mathbb{P}_{+}T^{*}M$ are Cartan geometries of type $(\mathrm{SL}(4,\mathbb{R}),P_{1,3}),$ where $P_{1,3}=P_{0}\ltimes P_{+}$ and $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R})$ is referred to as the structure group. At every $p\in\mathbb{P}_{+}T^{*}M,$ one finds that the almost complex structure
@@
is compatible since $J\in P_{0}.$ Clearly $P_{0}$ acts transitively on compatible almost complex structures on $\mathscr{F}_{p}$ and the stabilizer of $J$ is +×$\mathbb{R}^{+}\times\mathrm{CO}(2,\mathbb{R}).$ As a result, the fibers of the twistor bundle $\nu\colon\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ are
$\mathcal{T}^{[\nabla]}_{p}:=\nu^{-1}(p)=\mathrm{SL}(2,\mathbb{R})/\mathrm{SO}(2,\mathbb{R})\cong\mathbb{D}^{2}.$
@@
(3.11) $\mathcal{T}^{[\nabla]}\cong\mathcal{P}/P_{J}=\mathcal{P}\times_{P_{1,3}}P_{1,3}/P_{J}.$
Similarly to the two descriptions of $\mathcal{T}^{[g]}$ in § 2.4, one can give another description of $\mathcal{T}^{[\nabla]}$ using complexified Lagrangian null planes. To see this, by § 3.2, one knows that in the almost para-CR structure on $\mathbb{P}_{+}T^{*}M$ the subconformal structure $[g]\subset\mathrm{Sym}^{2}(\mathcal{F}^{*})$ and the conformal symplectic 2-form on the contact distribution $[\rho]\subset\bigwedge\nolimits^{2}\mathcal{F}^{*}$ are given by
@@
Allowing $[a_{1}:a_{2}]$ to be in $\mathbb{R}\mathbb{P}^{1}$ or $\mathbb{C}\mathbb{P}^{1},$ one obtains real and complex $\alpha$-planes, denoted by $\mathcal{A}^{\mathbb{R}}$ and $\mathcal{A}^{\mathbb{C}},$ respectively. The set of complex $\alpha$-planes without any real $\alpha$-plane i.e. $\mathcal{A}^{\mathbb{C}}_{p}\backslash\mathcal{A}^{\mathbb{R}}_{p},$ is isomorphic to $\mathbb{C}\mathbb{P}^{1}\backslash\mathbb{R}\mathbb{P}^{1}.$ Similarly to our discussion in § 2.4, after fixing an orientation, the fibers of $\mathcal{T}^{[\nabla]}$ can be identified with a connected component of $\mathcal{A}^{\mathbb{C}}_{p}\backslash\mathcal{A}^{\mathbb{R}}_{p}$ which is $\mathbb{D}^{2}.$ In other words, at every $p\in\mathbb{P}_{+}T^{*}M,$ using the action of $\mathrm{GL}(2,\mathbb{R})$ on $\mathscr{H}_{p}$ and $\mathscr{V}_{p},$ one can establish a bijection between elements of $\mathcal{A}_{p}^{\mathbb{C}}\backslash\mathcal{A}_{p}^{\mathbb{R}}$ and the dual $\mathrm{i}$-eigenspace for each compatible almost complex structure acting on $\mathscr{H}_{p}$ and $\mathscr{V}_{p}.$
@@
The 7-dimensional twistor bundle $\nu\colon\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ of an oriented projective structure []delimited-[][$\nabla$][ ∇ ] on a 3-manifold $M$ is equipped with a 2-nondegenerate CR structure of hypersurface type. The corank one distribution ^$\widehat{\mathscr{F}}\subset T\mathcal{T}^{[\nabla]}$ is the pre-image of the canonical contact distribution $\mathscr{F}\subset T\mathbb{P}_{+}T^{*}M,$ the Levi kernel $K\subset\widehat{\mathscr{F}}$ is the vertical distribution of the $\mathbb{D}^{2}$-bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ with its canonical complex structure $\mathscr{K}\oplus\overline{\mathscr{K}}=\mathbb{C}\otimes K$. With respect to the projection $\check{\pi}\colon\mathcal{P}\to\mathcal{T}^{[\nabla]}$ and Cartan connection (3.6), at each point $z\in\mathcal{T}^{[\nabla]},$ the holomorphic distribution ^$\widehat{\mathscr{H}}_{z}\subset\mathbb{C}\otimes\widehat{\mathscr{F}}_{z}$ is $\check{\pi}_{*}\mathrm{Ker}\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\}$ where
@@
^$\widehat{\mathscr{F}}=\nu_{*}^{-1}(\mathscr{F})=\check{\pi}_{*}\mathrm{Ker}\{\tau^{0}\}\subset T\mathcal{T}^{[\nabla]}.$
@@
(3.14) =$\mathcal{I}=\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\},$
where $\zeta^{i}$’s are as in (3.13). Using the structure equations (3.6) and (3.3), it is straightforward to show that $\mathcal{I}$ is integrable and defines a complex structure on $\widehat{\mathscr{F}}$ whose holomorphic distribution is given by ^=Ker^^$\widehat{\mathscr{H}}=\mathrm{Ker}\mathcal{I}\subset\mathbb{C}\otimes\widehat{\mathscr{F}}$.
@@
The complexification of degenerate directions of the Levi bracket are $\mathscr{K}\oplus\bar{\mathscr{K}}\subset\mathbb{C}\otimes K$ where $\mathscr{K}=\langle\check{\pi}_{*}\partial_{\zeta^{3}}\rangle.$ Using the symbol algebra of the almost para-CR structure one obtains
@@
Alternatively, at every $z\in\mathcal{T}^{[\nabla]},$ the 1-forms $\zeta^{1},\zeta^{2}$ in (3.13) can be viewed as the horizontal lift of the dual $\mathrm{i}$-eigenspace of the almost complex structure that corresponds to $z$ with respect to some almost para-CR Weyl structure. The 1-form $\zeta^{3}$ is the canonical holomorphic 1-form on $\mathrm{SL}(2,\mathbb{R})/\mathrm{SO}(2,\mathbb{R})$ that is invariant under the action of $\mathrm{SO}(2,\mathbb{R}).$ As before, this alternative construction of the CR structure is independent of the choice of the Weyl structure for the almost para-CR structure.
Remark 3.3.

### v3

@@ 3.2. Almost para-CR structure on the correspondence space
Given a projective structure $\nabla$] on a 3-manifold $M,$ the projectivized cotangent bundle, $\mathbb{P}T^{*}M,$ is equipped with an almost para-CR structure $\mathscr{F}=\mathscr{V}\oplus\mathscr{E}\subset T\mathbb{P}T^{*}M$ where $\mathscr{F}$ is the canonical contact distribution on $\mathbb{P}T^{*}M,$ $\mathscr{V}$ is the vertical distribution for the projection $\mathbb{P}T^{*}M\to M$ and $\mathscr{E}$ is obtained from the horizontal distributions defined by connections in $\nabla$].
@@
Recall that (2n1)(2n-1)-dimensional para-CR geometries are modeled by the flag variety of pairs of incident lines $\ell$ and hyperplanes $H$, i.e.
(3.5) $G/\penalty 50P_{1,n}\cong\mathbb{F}_{1,n}(\mathbb{R}^{n+1}):=\{(\ell,H)\ \vline\ \ell\subset H\subset\mathbb{R}^{n+1}\}\subset\mathbb{P}^{n}\times(\mathbb{P}^{n})^{*},$
where $G=\mathrm{SL}(n+1,\mathbb{R})$ and $P_{1,n}$ is the parabolic subgroup that stabilizes a chosen flag. The homogeneous space $\mathbb{F}_{1,n}(\mathbb{R}^{n+1})$ can be identified with $\mathbb{P}T^{*}\mathbb{P}^{n},$ i.e. the projectivized cotangent bundle of the projective space $\mathbb{P}^{n},$ equipped with its canonical contact structure and Legendrian foliations given by the fibers of the natural fibrations $\mathbb{P}T^{*}\mathbb{P}^{n}\to\mathbb{P}^{n}$ and $\mathbb{P}T^{*}\mathbb{P}^{n}\to(\mathbb{P}^{n})^{*}$.
Thus, almost para-CR structures in dimension 5 are Cartan geometries of type $(\mathrm{SL}(4,\mathbb{R}),P_{1,3})$ and are among the so-called regular and normal parabolic geometries, where $P_{1,3}=P_{0}\ltimes P_{+}$ in which $P_{+}$ is the 5-dimensional nilradical and $P_{0}=\mathbb{R}^{\times}\times\mathrm{GL}(2,\mathbb{R})\subset\mathrm{GL}(3,\mathbb{R}).$ In the case of oriented almost para-CR structures that correspond to oriented projective structures, one has $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R}).$ The harmonic curvature of 5-dimensional almost para-CR structures has three components, also known as the fundamental invariants, and can be written as
@@
The vanishing of $K_{H}$ implies local flatness, i.e. local equivalence to the flat model described previously. The vanishing of torsion components $T^{1}:\bigwedge\nolimits^{2}\mathscr{E}\to\mathscr{V}$ and $T^{2}:\bigwedge\nolimits^{2}\mathscr{V}\to\mathscr{E}$ is equivalent to the integrability of the rank 2 distributions $\mathscr{E}$ and $\mathscr{V},$ respectively. Moreover, the vanishing of $W$ is equivalent to the existence of a 4-parameter family of Legendrian surfaces that are null with respect to the indefinite subconformal structure on the contact distribution, e.g. see [KM24] for a review. An almost para-CR structure of dimension 2n12n-1 arises from a projective structure on an $n$-dimensional manifold in the manner described above, if and only if $W=0$ and $T^{2}=0,$ e.g. see [Čap05],
@@
(3.6) $\varphi=\begin{pmatrix}-\eta^{i}_{i}&\mu_{2}&\mu_{1}&\mu_{0}\\ \tau^{2}&\eta^{2}_{2}&\eta^{2}_{1}&\mu_{4}\\ \tau^{1}&\eta^{1}_{2}&\eta^{1}_{1}&\mu_{3}\\ \tau^{0}&\tau^{4}&\tau^{3}&\eta^{0}_{0}\end{pmatrix},\quad\Psi=\mathrm{d}\varphi+\varphi{\,{\wedge}\;}\varphi=\begin{pmatrix}0&\Gamma_{2}&\Gamma_{1}&\Gamma_{0}\\ 0&\Lambda^{2}_{2}&\Lambda^{2}_{1}&\Gamma_{4}\\ 0&\Lambda^{1}_{2}&\Lambda^{1}_{1}&\Gamma_{3}\\ 0&\Omega^{4}&\Omega^{3}&\Lambda^{0}_{0}\end{pmatrix}$,
@@
where, using (3.3) and working modulo $\{\tau^{0}\}$, the 2-forms entries $\Omega^{3}$ and $\Omega^{4}$ in $\Psi$ satisfy
(3.8) $\Omega^{3}\equiv W^{0}_{112}\tau^{1}{\,{\wedge}\;}\tau^{2},\quad\Omega^{4}\equiv W^{0}_{212}\tau^{1}{\,{\wedge}\;}\tau^{2}.$
@@
(3.9) $\mathscr{V}=\langle\partial_{p^{1}},\partial_{p^{2}}\rangle,\qquad\mathscr{E}=\langle v_{1},v_{2}\rangle,$
@@ 3.3. The twistor bundle and its 2-nondegenerate CR structure
Given an oriented projective structure $\nabla$] on $M,$ let $(\mathbb{P}_{+}T^{*}M;\mathscr{E},\mathscr{V})$ be its corresponding almost para-CR structure on the oriented projectivized cotangent bundle, $\mathbb{P}_{+}T^{*}M,$ with co-oriented contact distribution $\mathscr{F}=\mathscr{E}\oplus\mathscr{V}$. We define the twistor bundle of $\nabla$], denoted by $\nu\colon T^{[\nabla]}\to\mathbb{P}_{+}T^{*}M,$ as the bundle of almost complex structures on $\mathscr{F}$ that are compatible with its conformal symplectic 2-form and indefinite bilinear form. As a result, at every $p\in\mathbb{P}_{+}T^{*}M,$ any such compatible almost complex structure acts on $\mathscr{F}_{p}$ via the action of the structure group. Recall from § 3.2 that almost para-CR structures on $\mathbb{P}_{+}T^{*}M$ are Cartan geometries of type $(\mathrm{SL}(4,\mathbb{R}),P_{1,3}),$ where $P_{1,3}=P_{0}\ltimes P_{+}$ and $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R})$ is referred to as the structure group. At every $p\in\mathbb{P}_{+}T^{*}M,$ one finds that the almost complex structure
@@
is compatible since $J\in P_{0}.$ Clearly $P_{0}$ acts transitively on compatible almost complex structures on $\mathscr{F}_{p}$ and the stabilizer of $J$ is $\mathbb{R}^{+}\times\mathrm{CO}(2,\mathbb{R}).$ As a result, the fibers of the twistor bundle $\nu\colon\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ are
$\mathcal{T}^{[\nabla]}_{p}:=\nu^{-1}(p)=\mathrm{SL}(2,\mathbb{R})/\penalty 50\mathrm{SO}(2,\mathbb{R})\cong\mathbb{D}^{2}.$
@@
(3.11) $\mathcal{T}^{[\nabla]}\cong\mathcal{P}/\penalty 50P_{J}=\mathcal{P}\times_{P_{1,3}}P_{1,3}/\penalty 50P_{J}.$
Similarly to the two descriptions of $\mathcal{T}^{[g]}$ in § 2.4, one can give another description of $\mathcal{T}^{[\nabla]}$ using complexified Lagrangian null planes. To see this, by § 3.2, one knows that in the almost para-CR structure on $\mathbb{P}_{+}T^{*}M,$ the subconformal structure $g]\subset\mathrm{Sym}^{2}(\mathscr{F}^{*})$ and the conformal symplectic 2-form on the contact distribution $\rho]\subset\bigwedge\nolimits^{2}\mathscr{F}^{*}$ are given by
@@
Allowing $a_{1}:a_{2}]$ to be in $\mathbb{R}\mathbb{P}^{1}$ or $\mathbb{C}\mathbb{P}^{1},$ one obtains real and complex $\alpha$-planes at $p\in M$, denoted by $\mathcal{A}_{p}^{\mathbb{R}}$ and $\mathcal{A}_{p}^{\mathbb{C}},$ respectively. The set of complex $\alpha$-planes without any real $\alpha$-plane i.e. $\mathcal{A}^{\mathbb{C}}_{p}\backslash\mathcal{A}^{\mathbb{R}}_{p},$ is isomorphic to $\mathbb{C}\mathbb{P}^{1}\backslash\mathbb{R}\mathbb{P}^{1}.$ Similarly to our discussion in § 2.4, after fixing an orientation, the fibers of $\mathcal{T}^{[\nabla]}$ can be identified with a connected component of $\mathcal{A}^{\mathbb{C}}_{p}\backslash\mathcal{A}^{\mathbb{R}}_{p}$ which is $\mathbb{D}^{2}.$ In other words, at every $p\in\mathbb{P}_{+}T^{*}M,$ using the action of $\mathrm{GL}(2,\mathbb{R})$ on $\mathscr{E}_{p}$ and $\mathscr{V}_{p},$ one can establish a bijection between elements of $\mathcal{A}_{p}^{\mathbb{C}}\backslash\mathcal{A}_{p}^{\mathbb{R}}$ and the dual $\mathrm{i}$-eigenspace for each compatible almost complex structure acting on $\mathscr{E}_{p}$ and $\mathscr{V}_{p}.$
@@
The 7-dimensional twistor bundle $\nu\colon\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ of an oriented projective structure $\nabla$] on a 3-manifold $M$ is equipped with a 2-nondegenerate CR structure of hypersurface type. The corank one distribution $\widehat{\mathscr{F}}\subset T\mathcal{T}^{[\nabla]}$ is the pre-image of the canonical contact distribution $\mathscr{F}\subset T\mathbb{P}_{+}T^{*}M,$ the Levi kernel $K\subset\widehat{\mathscr{F}}$ is the vertical distribution of the $\mathbb{D}^{2}$-bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ with its canonical complex structure $\mathscr{K}\oplus\overline{\mathscr{K}}=\mathbb{C}\otimes K$. With respect to the projection $\check{\pi}\colon\mathcal{P}\to\mathcal{T}^{[\nabla]}$ and Cartan connection (3.6), at each point $z\in\mathcal{T}^{[\nabla]},$ the holomorphic distribution $\widehat{\mathscr{H}}_{z}\subset\mathbb{C}\otimes\widehat{\mathscr{F}}_{z}$ is $\check{\pi}_{*}\mathrm{Ker}\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\}$ where
@@
$\widehat{\mathscr{F}}=\nu_{*}^{-1}(\mathscr{F})=\check{\pi}_{*}\mathrm{Ker}\{\tau^{0}\}\subset T\mathcal{T}^{[\nabla]}.$
@@
(3.14) $\mathcal{I}=\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\},$
where $\zeta^{i}$’s are as in (3.13). Using the structure equations (3.6) and (3.3), it is straightforward to show that $\mathcal{I}$ is integrable and defines a complex structure on $\widehat{\mathscr{F}}$ whose holomorphic distribution is given by $\widehat{\mathscr{H}}=\mathrm{Ker}\mathcal{I}\subset\mathbb{C}\otimes\widehat{\mathscr{F}}$.
@@
The complexification of degenerate directions of the Levi bracket is $\mathscr{K}\oplus\bar{\mathscr{K}}\subset\mathbb{C}\otimes K$ where $\mathscr{K}=\langle\check{\pi}_{*}\partial_{\zeta^{3}}\rangle.$ Using the symbol algebra of the almost para-CR structure one obtains
@@
Alternatively, at every $z\in\mathcal{T}^{[\nabla]},$ the 1-forms $\zeta^{1},\zeta^{2}$ in (3.13) can be viewed as the horizontal lift of the dual $\mathrm{i}$-eigenspace of the almost complex structure that corresponds to $z$ with respect to some almost para-CR Weyl structure. The 1-form $\zeta^{3}$ is the canonical holomorphic 1-form on $\mathrm{SL}(2,\mathbb{R})/\penalty 50\mathrm{SO}(2,\mathbb{R})$ that is invariant under the action of $\mathrm{SO}(2,\mathbb{R}).$ As before, this alternative construction of the CR structure is independent of the choice of the Weyl structure for the almost para-CR structure.
Remark $3.$